OpenStudy (anonymous):

HELP PLEASE I NEED SOMEONE REALLY GOOD IN CALC 2 PLEASE HELP!

OpenStudy (anonymous):

what's your question? :)

OpenStudy (anonymous):

the Project problem

OpenStudy (anonymous):

you can use formula for length of the arc to answer your first question there (a): http://tutorial.math.lamar.edu/Classes/CalcII/ArcLength.aspx

OpenStudy (anonymous):

after you calculate the length, you can then use that length of the wire to calculate (b) using simple distance = time * speed equation

OpenStudy (anonymous):

and (c) is basically just replacing 'x' in the initial catenary equation with L/2 and unfolding it.

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

is this ok for letter a?

OpenStudy (anonymous):

dy/dx = (T/(ug)) sinh(ugx/T)(ug/T) = sinh(ugx/T) 1 + (dy/dx)^2 = 1 + sinh^2(ugx/T) = cosh^2(ugx/T) [1 + (dy/dx)^2]^(1/2)= cosh(ugx/T) integral of [1 + (dy/dx)^2]^(1/2) dx = integral of cosh(ugx/T) dx = [T/(ug)] sinh(ugx/T)